A patient is suspected of having low stomach acid, a condition known as hypochloridia. To determine whether the patient has this condition, her doctors take a 19.00 mL sample of her gastric juices and titrate the sample with 0.000300 M KOH. The gastric juice sample required 12.1 mL of the KOH titrant to neutralize it. Calculate the pH of the gastric juice sample. Assume the sample contained no ingested food or drink which might otherwise interfere with the titration.
Answer #1
Concepts and reason

The problem is based on the concept of the pH of an acid can be calculated by titrating it against a base of known concentration as the gastric juice in human stomach is acidic in nature because it contains HCl{\rm{HCl}} .

Fundamentals

The principle of titration is that number of moles of acid (nA)\left( {{n_{\rm{A}}}} \right) react with equal number of moles (nB)\left( {{n_{\rm{B}}}} \right) of base.

nA=nB{n_{\rm{A}}} = {n_{\rm{B}}} …… (1)

The pH of an acidic solution is

pH=log[H+]{\rm{pH}} = - \log \left[ {{{\rm{H}}^ + }} \right] …… (2)

Here, [H+]\left[ {{{\rm{H}}^ + }} \right] is concentration of hydrogen ion.

Molarity (M) is defined as the number of moles of solute distributed in per litre of a solution.

M=n(moles)V(L)M = {\rm{ }}\frac{{n\left( {{\rm{moles}}} \right)}}{{V\left( {\rm{L}} \right)}} …… (3)

Here nn is the number of moles and VV is the volume of the solution in liters.

Number of moles (n)\left( n \right) in a solution is the product of molarity (M) and volume of the solution (V). Rearranging equation (3) this equation is obtained.

n=M×Vn = M \times V …… (4)

Molarity of KOH{\rm{KOH}} solution (MNaOH)\left( {{M_{{\rm{NaOH}}}}} \right) is 0.000300M{\rm{0}}{\rm{.000300 M}} and the volume of KOH{\rm{KOH}} used (VKOH)\left( {{V_{{\rm{KOH}}}}} \right) is 12.1 mL

Number of moles of KOH (nKOH)({n_{KOH}}) can be determined from equation (2).

Substitute, 0.000300M0.000300{\rm{ M}} for MM and 12.1mL12.1{\rm{ mL}} for VV in equation (2).

n=0.000300M×12.1mL=(0.000300molL×12.1mL×1L1000mL)=3.63×106mol\begin{array}{c}\\n = 0.000300{\rm{ M}} \times 12.1{\rm{ mL}}\\\\ = \left( {0.000300{\rm{ }}\frac{{{\rm{mol}}}}{{\rm{L}}} \times 12.1{\rm{ mL}} \times \frac{{1{\rm{ L}}}}{{1000{\rm{ mL}}}}} \right)\\\\ = 3.63 \times {10^{ - 6}}{\rm{ mol}}\\\end{array}

Number of moles of HCl{\rm{HCl}} (nHCl)\left( {{n_{{\rm{HCl}}}}} \right) will be equal to the number of moles of KOH{\rm{KOH}} (nKOH)\left( {{n_{{\rm{KOH}}}}} \right)

nHCl=nKOH{n_{{\rm{HCl}}}} = {n_{{\rm{KOH}}}}

Substitute the value of nKOH{n_{{\rm{KOH}}}} in the above expression.

nHCl=3.63×106mol{n_{{\rm{HCl}}}} = 3.63 \times {10^{ - 6}}{\rm{ mol}}

Molarity of HCl{\rm{HCl}} (nHCl)\left( {{n_{{\rm{HCl}}}}} \right) can be calculated from equation (3).

MHCl=nHClVHCl{M_{{\rm{HCl}}}} = \frac{{{n_{{\rm{HCl}}}}}}{{{V_{{\rm{HCl}}}}}}

Substitute 3.63×106mol3.63 \times {10^{ - 6}}{\rm{ mol}} for nHCl{n_{{\rm{HCl}}}} and 19.00 mL for VHCl{V_{{\rm{HCl}}}} to the above expression.

MHCl=3.63×106mol19.00mL=3.63×106mol19.00mL×1000mL1L=1.91×104molL1\begin{array}{c}\\{M_{{\rm{HCl}}}} = \frac{{3.63 \times {{10}^{ - 6}}{\rm{ mol}}}}{{19.00{\rm{ mL}}}}\\\\ = \frac{{3.63 \times {{10}^{ - 6}}{\rm{ mol}}}}{{19.00{\rm{ mL}}}} \times \frac{{1000{\rm{ mL}}}}{{1{\rm{ L}}}}\\\\ = 1.91 \times {10^{ - 4}}{\rm{ mol }}{{\rm{L}}^{ - 1}}\\\end{array}

Concentration of HCl{\rm{HCl}} (MHCl)\left( {{M_{{\rm{HCl}}}}} \right) is equal to concentration of H+{{\rm{H}}^{\rm{ + }}} ions.

[H+]=[HCl][H+]=MHCl\begin{array}{c}\\\left[ {{{\rm{H}}^ + }} \right] = \left[ {{\rm{HCl}}} \right]\\\\\left[ {{{\rm{H}}^ + }} \right] = {M_{{\rm{HCl}}}}\\\end{array}

Substitute 1.91×104molL11.91 \times {10^{ - 4}}{\rm{ mol }}{{\rm{L}}^{ - 1}} for MHCl{M_{{\rm{HCl}}}} .

[H+]=1.91×104molL1\left[ {{{\rm{H}}^ + }} \right] = 1.91 \times {10^{ - 4}}{\rm{ mol }}{{\rm{L}}^{ - 1}}

The pH of the solution can be calculated from the equation (2).

Substitute, 1.91×104molL11.91 \times {10^{ - 4}}{\rm{ mol }}{{\rm{L}}^{ - 1}} for [H+]\left[ {{{\rm{H}}^{\rm{ + }}}} \right] in equation (2).

pH=log[1.91×104molL1]=3.72\begin{array}{c}\\{\rm{pH}} = - \log \left[ {1.91 \times {{10}^{ - 4}}{\rm{ mol }}{{\rm{L}}^{ - 1}}} \right]\\\\ = 3.72\\\end{array}

Ans:

The pH of the gastric juice is 3.723.72 .